OpenStudy (charlotte123):

For the following system, if you isolated y in the first equation to use the Substitution Method, what expression would you substitute into the second equation?: 2x + y = 8 −x − y = −5 8 + 2x 8 − 2x −8 − 2x −8 + 2x

OpenStudy (tkhunny):

Well, do it. Isolate y in the first equation.

OpenStudy (charlotte123):

@tkhunny It's a pretest, so I don't know how to do that yet D:

OpenStudy (anonymous):

2x + y = 8 You want to get y by itself, so subtract 2x from both sides 2x - 2x + y = 8 - 2x (combine like terms) y = 8 - 2x Do you understand ?

OpenStudy (charlotte123):

@kelliegirl33 that was the answer I put as my choice! :D I just didn't know I did it correctly XD thanks kellie ^-^

OpenStudy (charlotte123):

tkhunny ya wanna add something? :)

OpenStudy (anonymous):

no problem...glad to help :)

OpenStudy (tkhunny):

Perfect. Then you should not do it. If it's a pretest, would it make any sense at all for you get smoke very question? You should learn it throughout the course of your semester. In this case, subtraction solves the problem, as demonstrated by kelliegirl33.

OpenStudy (anonymous):

I believe she was just double checking her answer

OpenStudy (charlotte123):

@tkhunny but now, thanks to kellie, I know how to do it ^-^ this way when I read the lesson, I can already know how to solve problems like this one -.-

OpenStudy (charlotte123):

@kelliegirl33 ya know I wanted to send u a message a long time ago, but now ur here, I just want to say thanks for help :) not only did ya just help me now, but let me just say that I have passed my history tests all thanks to ur help ^-^ lol if there was more than 1 medal I could give, I would XD

OpenStudy (tkhunny):

Therein lies the problem. The rules of engagement are to SHOW YOUR WORK. @charlotte123 , no matter what you are doing, please show YOUR work up front. Don't make someone else do it when you already have done it. Plus, it doesn't quite make sense that you said you did not know how to do it.

OpenStudy (anonymous):

thank you...I try to help when I can

OpenStudy (charlotte123):

@tkhunny thanks for the help :) you have a great day ^-^

OpenStudy (charlotte123):

@kelliegirl33 haha XD np :) thanks again! :D

OpenStudy (anonymous):

no problem :)